# Improving Underwater Acoustic Classification Through Learnable Gabor Filter Convolution and Attention Mechanisms

### v1

@@ 4.1 Datasets
In the context of radiated ship noise, the DeepShip database (irfan2021deepship) is widely used. It is derived from the Ocean Networks Canada Society database, [^1] https://oceannetworks.ca which contains unlabelled underwater acoustic data. DeepShip recordings were captured between May 2016 and October 2018 at the Strait of Georgia delta node and are segmented into three time windows with hydrophones positioned at depths of 141 m, 147 m, and 144 m. Data acquisition was performed using an icListen smart hydrophone offering a frequency range of 1 Hz to 12 kHz and a sampling rate of 32,000 Hz. Vessel labels were assigned based on Automatic Identification System (AIS) data and include four categories: tugboat, passenger ship, cargo ship, and oil tanker. To ensure clean samples, only recordings containing signals from a single vessel within a 2 km radius were retained. These recordings were truncated and discarded once a second ship entered this range. The final dataset consisted of 613 recordings from 265 distinct vessels, totalling approximately 47 hours. It also provides information on the vessel’s distance from the hydrophone at three different time points to track ship trajectories. While (irfan2021deepship) proposes a five-class scheme including background noise, only the four vessel categories are present in the dataset. Some studies address this by sourcing background noise recordings from other datasets. In this study, however, only the four provided classes are used, without the inclusion of background noise recordings.
@@ 4.2 Classification Metrics
In a classification pipeline, model performance is typically evaluated based on overall accuracy, defined as the ratio of correct classifications out of the total number of instances. However, accuracy alone can lead to misinterpretation of the results, since it does not distinguish between true positives ($tp$), true negatives ($tn$), false positives ($fp$) and false negatives ($fn$). Metrics such as Precision ($\frac{tp}{tp+fp}$), Recall ($\frac{tp}{tp+fn}$), and the F1-score, which is the weighted harmonic mean of precision and recall, usually provide a global intuition about the results. However, the Matthews Correlation Coefficient (MCC) (matthews1975comparison) is usually a more appropriate metric (domingos2022survey), as it assigns a high score to predictions that are correct for all four base categories ($tp$, fpfp, tntn and $fp$). This coefficient is defined by the following expression:
@@
While most studies in the literature use micro-averaged accuracy to report their conclusions, there are several limitations to relying on accuracy alone. For example, in this four-class classification problem, a random predictor would achieve approximately 25% accuracy, whereas the same predictor would yield a MCC of 0%, correctly indicating randomness. To minimise the impact of data imbalance on the results, MCC is used as the reference metric.
@@ 4.3 Framework and Hyper-parameters
To encompass the most common classification strategies, the performance of the GSE ResNeXt model was compared with that of several well-established models in the literature: a classic residual network (ResNet) (he2016deep), serving as a baseline to assess the improvements introduced in this study over the traditional architecture; MobileNetV2 (sandler2018mobilenetv2), an efficient lightweight CNN that employs inverted residuals with linear bottlenecks, enabling analysis of the relationship between model compactness and accuracy; and Xception (chollet2017xception), an architecture that employs multi-scale residual blocks and depthwise separable convolutions, allowing comparison with the multi-path feature extraction strategy.
To emphasise the lower-frequency components of ship-radiated noise, Constant-Q Transform (CQT) spectrograms (brown1991calculation) were selected as input, following conclusions drawn in recent literature (domingos2022investigation; irfan2021deepship). The CQT enhances low-frequency resolution using geometrically spaced filters and a constant quality factor, making it well-suited for identifying harmonic structures in underwater acoustic signals. The CQT hyperparameters were selected based on the characteristics of the audio signals in the DeepShip dataset, as shown in Table 2. Audio segments of 30 seconds in duration were used, with the SpecAugment strategy applied as a data augmentation step to improve training robustness, in line with recent studies (tian2023joint; hong2021underwater; liu2021underwater; zhou2023attentionbased). Training was conducted using the AdamW optimiser with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$ for regularisation, over 50 epochs with a mini-batch size of 64.
@@
As Task 2 preserves the temporal consistency of the audio files, a more complete set of test configurations was established in the analysis. Each audio file was divided into five equal-length segments, each of which was included in a different fold. After this division, they were segmented into 30-second pieces for spectrogram generation. Thus, Fold 1 contains the beginning of each audio file, Fold 5 contains the end, and Folds 2, 3, and 4 contain the intermediate parts. For a complete analysis of the results, two folds were considered for training: the initial segment of the audio (Fold 1) and the ending (Fold 5). The results were then analysed using the remaining segments as the testing subset, producing only two combinations: training on the first fold and testing on the last four, and training on the last fold and testing on the first four. As at least five 30-second segments were needed for Task 2, and to maintain consistency with other tasks, only audio files with a duration of more than one minute and 30 seconds were considered for all the three tasks.
To evaluate the model’s ability to separate classes across all three tasks, an analysis based on latent features was conducted. The feature representations obtained before the classification layer were extracted using the test subset and projected into two dimensions using the UMAP algorithm (mcinnes2020umap). A comparison was made between the latent space representations categorised by the true labels in the dataset and those categorised according to the model’s predictions. These outcomes provide insight into the model’s capacity to extract meaningful features from the training data and generalise effectively to the test data. It also gives insight on the isolation between training and test subsets.
@@ 5 Results and Discussion
Model Parameters FLOPs MCC (%)
Task 1 Task 2 Task 3
MobileNet V2 2.2M 0.4G 99.19 ± 0.15 65.12 ± 2.35 59.08 ± 6.22
ResNet 18 11.2M 2.1G 99.42 ± 0.33 75.34 ± 2.34 64.15 ± 3.64
Xception 20.8M 5.6G 99.63 ± 0.14 76.55 ± 0.89 63.00 ± 3.29
GSE ResNeXt 14.7M 3.0G 99.53 ± 0.15 78.53 ± 2.50 64.86 ± 4.83
Table 3: Baseline results obtained for the three classification tasks.
Analysis of the results presented in Table 3 shows that Task 1 achieves an MCC of around 99% for all four models, showing that the similarity between adjacent audio segments is sufficient for the model to extract almost 100% of the classification information from the training dataset. Variations in model size and architecture do not interfere with the final classification results, as the models are probably only fitting the training data and not providing enough information to evaluate their application to unseen data. While these results demonstrate that the input data can be classified using representative data, they also highlight the need for a more comprehensive evaluation of the approach, focusing on tasks 2 and 3.
The results obtained for Task 3 show an accuracy drop of approximately 35 MCC percentage points compared to Task 1, reflecting the improved isolation between the training and test subsets. With reduced similarities between subsets, since the same ship recording was not present in both, the classification task becomes more complex, resulting in a significant decrease in performance. This task also exhibited greater variability in performance across different folds, with a standard deviation exceeding six percentage points when using MobileNet V2, the worst-performing model. This variation raises concerns regarding the similarities between different ship classes and their taxonomy, since classification results differ significantly depending on the folds division, as illustrated in Figure 2. Since there are different recordings in each fold, radiated noise of the same ship is unlikely to be placed on both the training and test datasets, making the accurate classification of such signals challenging.
@@
Task 3 results also suggest that the taxonomy of classes based on ship types is unrepresentative, as the classes may not be clearly separated. To evaluate this hypothesis, the analysis based on latent features was performed, as shown in Figure 3. The top row shows the latent space representations categorised by the true labels in the dataset, while the bottom row shows the same representations categorised according to the model’s predictions. The difference between the tasks is noticeable, with the class boundaries in Task 1 being extremely well-defined, whereas those for Task 3 are mixed into a complex shape. Since Task 1 includes similar portions of the same recording in both the training and test datasets, the separation of the features is well-defined, producing a model that can distinguish between classes based on their training instances. Task 3, on the other hand, evaluates the inverse relationship, due to the complete separation of entire ship recordings it focuses on model performance in the absence of data leakage between the training and test sets. The complex boundaries obtained for the test dataset reflect this characteristic. Although the features were separated during training, those extracted from different classes were similar enough to be placed in sometimes overlapping coordinates in the latent space during testing. The conclusions from this analysis point in two directions. First, the training subset alone cannot represent the entire distribution of unseen data, meaning that even if the model achieves 100% accuracy on the training set, it will not generalise well enough to data outside that distribution. Second, there is no ideal way to categorise ship recordings into clearly defined classes, which tend to overlap in practice (e.g., a cargo ship and a tanker may sound nearly identical). A low-level taxonomy focusing on the physical characteristics of ships may offer a better solution in this case.
@@
Figure 3: The two-dimensional visualisation of the features obtained after the feature extraction stage for the three tasks considered. The top row displays the ground truth labels, while the bottom row shows the predicted labels. While the latent space for Task 1 exhibits clearer separation between classes, Task 3 reveals a more complex structure, even in the ground truth labels. Task 2 presents a marginally separable space, with minor overlaps between classes.
@@
The analysis of Task 2 aims to address the open discussions about the lack of distinctiveness between the group classifications from Task 3. Since the training and test datasets contain the same recordings, the taxonomy of classes can be better explored. Additionally, using consecutive audio segments, rather than the overlapping ones from Task 1, mitigates the data leakage problem. The results presented in Table 3 show the performance of the classifiers for Task 2, which yielded accuracy levels between those observed in Tasks 1 and 3. The latent space representations obtained for this task are also intermediate between those of Tasks 1 and 3, exhibiting a moderate level of class separation with some overlap at the boundaries, while still maintaining a complex overall structure. The middle region of the latent space shows intersection of classes, suggesting the similar conclusions obtained for Task 3 about the taxonomy. The smaller standard deviation between folds also suggest that there is a smaller impact of the folds division in the final result. It points out that when using the same recordings, thereby the same ships, for training and test, the results are more consistent, therefore the distributions between the subsets are similar even without data leakage. This consistency is fundamental to compare the results obtained on different pipelines, alleviating dataset biases in the classification results.
@@
Considering the Task 2 results, the comparison between the baseline models suggest that exploring not only the depth but also the width of the networks, i.e. by adding parallel operations to their branches, improves performance. In this context, the GSE ResNeXt model achieved a 2.58% higher MCC than the Xception model, which obtained the second-best result. This demonstrates that combining attention and Gabor convolutional layers is also beneficial for this classification problem. As the Gabor convolutional layer, which is the first layer in the GSE ResNeXt architecture, approximates a two-dimensional band-pass filter, the features obtained after this layer represent the various two-dimensional frequency bands present in the input spectrogram. Applying channel attention after these features produces a frequency selection behaviour, focusing the training towards the relevant spatial frequencies in the input representation, thereby improving noise robustness and convergence.
@@
To evaluate the noise robustness of the Gabor convolutional layers, an analysis was conducted on the initial kernels learned by ResNeXt and GSE ResNeXt models. As shown in Figure 4, the kernels learned using traditional convolutions converged to a sparser representation with greater variation in magnitude between neighbouring pixels. Conversely, Gabor convolutional kernels naturally preserve spatial frequencies, contributing to a improved noise robustness. The convergence of Gabor kernels to a diverse range of natural frequencies and angles contributed to a more comprehensive representation of the input data, which also influenced the final accuracy.
@@
To evaluate the training convergence, the epoch at which the training achieved a steady state was identified. In this experiment, the steady state was defined as the epoch at which the validation MCC was within 2% of the final value for three consecutive epochs. Using Gabor convolutions in the first layer, the steady state was achieved at epoch 33; however, it was only achieved at epoch 46 when traditional convolutions were used. This represents a 28% improvement, as can be seen in Figure 5. The Gabor convolution layer significantly improved the convergence of the ResNeXt model, also improving the final performance. It is important to note that the improvements observed for Task 2 were also observed for Task 3, as demonstrated by the superior classification performance of the GSE ResNeXt model. Nothing can be said about Task 1, since the classification accuracy was almost perfect in all baseline models.
Figure 5: Comparison of the MCC values obtained on the validation dataset during the training stage of a ResNeXt model with traditional convolutions and a ResNeXt model with Gabor convolutions in the first layer. The red dashed line indicates the epoch at which training reached a steady state. The use of Gabor convolutions resulted in a more stable validation metric and faster convergence compared to traditional convolutions.

### v2

@@ 4.1 Datasets
In the context of radiated ship noise, the DeepShip database [23] is widely used. It is derived from the Ocean Networks Canada Society database, [^1] https://oceannetworks.ca which contains unlabelled underwater acoustic data. DeepShip recordings were captured between May 2016 and October 2018 at the Strait of Georgia delta node and are segmented into three time windows with hydrophones positioned at depths of 141 m, 147 m, and 144 m. Data acquisition was performed using an icListen smart hydrophone offering a frequency range of 1 Hz to 12 kHz and a sampling rate of 32,000 Hz. Vessel labels were assigned based on Automatic Identification System (AIS) data and include four categories: tugboat, passenger ship, cargo ship, and oil tanker. To ensure clean samples, only recordings containing signals from a single vessel within a 2 km radius were retained. These recordings were truncated and discarded once a second ship entered this range. The final dataset consisted of 613 recordings from 265 distinct vessels, totalling approximately 47 hours. It also provides information on the vessel’s distance from the hydrophone at three different time points to track ship trajectories. While [23] proposes a five-class scheme including background noise, only the four vessel categories are present in the dataset. Some studies address this by sourcing background noise recordings from other datasets. In this study, however, only the four provided classes are used, without the inclusion of background noise recordings.
@@ 4.2 Classification Metrics
In a classification pipeline, model performance is typically evaluated based on overall accuracy, defined as the ratio of correct classifications out of the total number of instances. However, accuracy alone can lead to misinterpretation of the results, since it does not distinguish between true positives ($tp$), true negatives ($tn$), false positives ($fp$) and false negatives ($fn$). Metrics such as Precision ($\frac{tp}{tp+fp}$), Recall ($\frac{tp}{tp+fn}$), and the F1-score, which is the weighted harmonic mean of precision and recall, usually provide a global intuition about the results. However, the Matthews Correlation Coefficient (MCC) [31] is usually a more appropriate metric [11], as it assigns a high score to predictions that are correct for all four base categories ($tp$, fpfp, tntn and $fp$). This coefficient is defined by the following expression:
@@
While most studies in the literature use micro-averaged accuracy to report their conclusions, there are several limitations to relying on accuracy alone. For example, in this four-class classification problem, a random predictor would achieve approximately 25% accuracy, whereas the same predictor would yield a MCC of 0%, correctly indicating randomness. To minimise the impact of data imbalance on the results, MCC is used as the reference metric, complementing the accuracy information.
@@ 4.3 Framework and Hyper-parameters
To encompass the most common classification strategies, the performance of the GSE ResNeXt model was compared with that of several well-established models in the literature: a classic residual network (ResNet) [19], serving as a baseline to assess the improvements introduced in this study over the traditional architecture; MobileNetV2 [40], an efficient lightweight CNN that employs inverted residuals with linear bottlenecks, enabling analysis of the relationship between model compactness and accuracy; Xception [9], an architecture that employs multi-scale residual blocks and depthwise separable convolutions, allowing comparison with the multi-path feature extraction strategy; MobileViT [34], which combines convolutional layers with vision transformers to balance local and global feature representations; and AST (Audio Spectrogram Transformer) [16], a transformer-based model designed for audio spectrogram inputs, enabling evaluation of attention-driven architectures in underwater acoustic classification.
To emphasise the lower-frequency components of ship-radiated noise, Constant Q Transform (CQT) spectrograms [5] were selected as input, following conclusions drawn in recent literature [12, 23]. The CQT enhances low-frequency resolution using geometrically spaced filters and a constant quality factor, making it well-suited for identifying harmonic structures in underwater acoustic signals. The CQT hyperparameters were selected based on the characteristics of the audio signals in the DeepShip dataset, as shown in Table 2. The choice of using a sample rate of 32,000 samples per second ensures that the full spectral content of the signal is preserved. This allows the feature extraction process to select relevant information during training rather than constraining the input to a predefined range. Audio segments of 30 seconds in duration were used, with the SpecAugment [36] strategy applied as a data augmentation step to improve training robustness, in line with recent studies [43, 21, 29, 60]. Training was conducted using the AdamW optimiser with a learning rate of $1\times 10^{-3}$ and a weight decay of $1\times 10^{-4}$ for regularisation, over 50 epochs with a mini-batch size of 64. Cross-entropy loss with label smoothing was used as criterion during the training procedure.
@@
As Task 2 preserves the temporal consistency of the audio files, a more complete set of test configurations was established in the analysis. Each audio file was divided into five equal-length segments, each of which was included in a different fold. After this division, they were segmented into 30-second pieces for spectrogram generation. Thus, Fold 1 contains the beginning of each audio file, Fold 5 contains the end, and Folds 2, 3, and 4 contain the intermediate parts. For a complete analysis of the results, two folds were considered for training: the initial segment of the audio (Fold 1) and the ending (Fold 5). The results were then analysed using the remaining segments as the testing subset, producing only two combinations: training on the first fold and testing on the last four, and training on the last fold and testing on the first four. As at least five 30-second segments were needed for Task 2, and to maintain consistency with other tasks, only audio files with a duration of more than one minute and 30 seconds were considered for all the three tasks. The total number of 30-second audio segments per class and their respective proportions in the dataset are shown in Table 3, which also indicates that class balance was maintained.
Tug Tanker Cargo Passenger Ship
Number of Segments 2668 2647 2519 2905
Proportion (%) 24.85% 24.64% 23.46% 27.06%
Table 3: The dataset class distribution including the number of 30 second segments and its proportion of total dataset.
To evaluate the model’s ability to separate classes across all three tasks, an analysis based on latent features was conducted. The feature representations obtained before the classification layer were extracted using the test subset and projected into two dimensions using the UMAP algorithm [32]. A comparison was made between the latent space representations categorised by the true labels in the dataset and those categorised according to the model’s predictions. These outcomes provide insight into the model’s capacity to extract meaningful features from the training data and generalise effectively to the test data. It also gives insight on the isolation between training and test subsets.
@@ 5 Results and Discussion
Model Params FLOP MCC (%) Accuracy (%)
Task 1 Task 2 Task 3 Task 1 Task 2 Task 3
MobileNet V2 2.2M 0.4G 99.19 $\pm$ 0.15 65.12 $\pm$ 2.35 59.08 $\pm$ 6.22 99.39 $\pm$ 0.11 73.82 $\pm$ 1.78 69.01 $\pm$ 4.93
ResNet 18 11.2M 2.1G 99.42 $\pm$ 0.33 75.34 $\pm$ 2.34 64.15 $\pm$ 3.64 99.56 $\pm$ 0.25 81.46 $\pm$ 1.79 72.75 $\pm$ 3.06
Xception 20.8M 5.6G 99.63 $\pm$ 0.14 76.55 $\pm$ 0.89 63.00 $\pm$ 3.29 99.72 $\pm$ 0.10 82.39 $\pm$ 0.66 72.03 $\pm$ 2.72
AST 5.9M 7.8G 92.61 $\pm$ 1.42 59.33 $\pm$ 0.53 55.05 $\pm$ 6.01 94.45 $\pm$ 1.07 69.48 $\pm$ 0.43 66.10 $\pm$ 4.71
MobileViT 5.0M 1.7G 98.65 $\pm$ 0.13 65.51 $\pm$ 0.71 57.10 $\pm$ 5.78 98.99 $\pm$ 0.09 74.05 $\pm$ 0.52 67.66 $\pm$ 4.48
GSE ResNeXt 14.7M 3.0G 99.53 $\pm$ 0.15 78.53 $\pm$ 2.50 64.86 $\pm$ 4.83 99.65 $\pm$ 0.11 83.85 $\pm$ 1.85 73.31 $\pm$ 3.75
Table 4: Baseline results obtained for the three classification tasks. Data is provided as mean $\pm$ standard deviation across folds, and accuracy is presented as complementary information alongside MCC.
Analysis of the results presented in Table 4 shows that Task 1 achieves an MCC of around 99% for four out of the six models, showing that the similarity between adjacent audio segments is sufficient for the model to extract almost 100% of the classification information from the training dataset. Variations in model size and architecture have minor interference with the final classification results, as the models are probably only fitting the training data and not providing enough information to evaluate their application to unseen data. While these results demonstrate that the input data can be classified using representative data, they also highlight the need for a more comprehensive evaluation of the approach, focusing on tasks 2 and 3.
The results obtained for Task 3 show an accuracy drop of approximately 35 MCC percentage points compared to Task 1, reflecting the improved isolation between the training and test subsets. With reduced similarities between subsets, since the same ship recording was not present in both, the classification task becomes more complex, resulting in a significant decrease in performance. This task also exhibited greater variability in performance across different folds, with a standard deviation exceeding six percentage points when using MobileNet V2 and AST, two of the worst-performing models. This variation raises concerns regarding the similarities between different ship classes and their taxonomy, since classification results differ significantly depending on the folds division, as illustrated in Figure 2. Since there are different recordings in each fold, radiated noise of the same ship is unlikely to be placed on both the training and test datasets, making the accurate classification of such signals challenging.
@@
Task 3 results also suggest that the taxonomy of classes based on ship types is unrepresentative, as the classes may not be clearly separated. To evaluate this hypothesis, the analysis based on latent features was performed using the GSE ResNeXt model, as shown in Figure 3. The top row shows the latent space representations categorised by the true labels in the dataset, while the bottom row shows the same representations categorised according to the model’s predictions. The difference between the tasks is noticeable, with the class boundaries in Task 1 being extremely well-defined, whereas those for Task 3 are mixed into a complex shape. Since Task 1 includes similar portions of the same recording in both the training and test datasets, the separation of the features is well-defined, producing a model that can distinguish between classes based on their training instances. Task 3, on the other hand, evaluates the inverse relationship, due to the complete separation of entire ship recordings it focuses on model performance in the absence of data leakage between the training and test sets. The complex boundaries obtained for the test dataset reflect this characteristic. Although the features were separated during training, those extracted from different classes were similar enough to be placed in sometimes overlapping coordinates in the latent space during testing. The conclusions from this analysis point in two directions. First, the training subset alone cannot represent the entire distribution of unseen data, meaning that even if the model achieves 100% accuracy on the training set, it will not generalise well enough to data outside that distribution. Second, there is no ideal way to categorise ship recordings into clearly defined classes, which tend to overlap in practice (e.g., a cargo ship and a tanker may sound nearly identical). A low-level taxonomy focusing on the physical characteristics of ships may offer a better solution in this case.
@@
Figure 3: The two-dimensional visualisation of the features obtained after the feature extraction stage in the GSE ResNeXt model for the three tasks considered. The top row displays the ground truth labels, while the bottom row shows the predicted labels. While the latent space for Task 1 exhibits clearer separation between classes, Task 3 reveals a more complex structure, even in the ground truth labels. Task 2 presents a marginally separable space, with minor overlaps between classes.
@@
The analysis of Task 2 aims to address the open discussions about the lack of distinctiveness between the group classifications from Task 3. Since the training and test datasets contain the same recordings, the taxonomy of classes can be better explored. Additionally, using consecutive audio segments, rather than the overlapping ones from Task 1, mitigates the data leakage problem. The classification results in Table 4 are complemented by the confusion matrix for Task 2 using the GSE ResNeXt model in Figure 4, which highlights the distribution of correct and misclassified samples across classes. Task 2 achieved accuracy levels between those observed in Tasks 1 and 3. The latent space representations obtained for this task are also intermediate between those of Tasks 1 and 3, exhibiting a moderate level of class separation with some overlap at the boundaries, while still maintaining a complex overall structure. The middle region of the latent space shows intersection of classes, suggesting the similar conclusions obtained for Task 3 about the taxonomy. The smaller standard deviation between folds also suggest that there is a smaller impact of the folds division in the final result. It points out that when using the same recordings, thereby the same ships, for training and test, the results are more consistent, therefore the distributions between the subsets are similar even without data leakage. This consistency is fundamental to compare the results obtained on different pipelines, alleviating dataset biases in the classification results.
Figure 4: Confusion matrix obtained for Task 2 using the GSE ResNeXt model. The matrix shows the distribution of correct predictions and misclassifications across ship classes.
@@
Considering the Task 2 results, the comparison between the baseline models suggest that exploring not only the depth but also the width of the networks, i.e. by adding parallel operations to their branches, improves performance. In this context, the GSE ResNeXt model achieved a 2.58% higher MCC than the Xception model, which obtained the second-best result. This demonstrates that combining attention and Gabor convolutional layers is also beneficial for this classification problem. The Gabor convolutional layer, the first layer in the GSE ResNeXt architecture, approximates a two-dimensional band-pass filter. Consequently, the features extracted after this layer represent the various two-dimensional frequency bands present in the input spectrogram, as illustrated in Figure 5, which shows the feature maps obtained for an example spectrogram. It is possible to notice how these maps exhibit diverse textures across channels. Applying channel attention after these features introduces a frequency-selection behaviour, focusing the training towards the relevant spatial frequencies in the input representation and thereby enhancing feature discrimination, noise robustness, and convergence.
Figure 5: An example of input spectrogram (a) and the 12 feature maps with the highest activation peaks extracted from the Gabor convolution layer of the GSE ResNeXt model (b). Maps are ordered by activation magnitude, illustrating the most prominent spatial features learned for the given input.
@@
To evaluate the noise robustness of the Gabor convolutional layers, an analysis was conducted on the initial kernels learned by ResNeXt and GSE ResNeXt models. As shown in Figure 6, the kernels learned using traditional convolutions converged to a sparser representation with greater variation in magnitude between neighbouring pixels. Conversely, Gabor convolutional kernels naturally preserve spatial frequencies, contributing to a improved noise robustness. The convergence of Gabor kernels to a diverse range of natural frequencies and orientations contributed to a more comprehensive representation of the input data, which also influenced the final classification accuracy.
@@
To evaluate the training convergence, the epoch at which the training achieved a steady state was identified. In this experiment, the steady state was defined as the epoch at which the validation MCC was within 2% of the final value for three consecutive epochs. Using Gabor convolutions in the first layer, the steady state was achieved at epoch 33 for Task 2; however, it was only achieved at epoch 46 when traditional convolutions were used. This represents a 28% improvement, as can be seen in Figure 7. The Gabor convolution layer significantly improved the convergence of the ResNeXt model, also improving the final performance. It is important to note that the improvements observed for Task 2 were also observed for Task 3, where the final state was reached at epoch 16 with Gabor convolutions compared to 43 with traditional ones, an improvement of 62%, further supported by the superior classification performance of the GSE ResNeXt model. Nothing can be said about Task 1, since the classification accuracy was almost perfect in all baseline models.
Figure 7: Comparison of the MCC values obtained on the validation dataset during the training stage of a SE ResNeXt model with traditional convolutions and a SE ResNeXt model with Gabor convolutions in the first layer. The red dashed line indicates the epoch at which training reached a steady state. The use of Gabor convolutions resulted in a more stable validation metric and faster convergence compared to traditional convolutions.